A pure sample of 226Ra contains 2.6 × 1013 atoms of the isotope. If the half-life of 226Ra = 1.6 × 103 years, what is the decay rate of this sample? (1 Ci = 3.7 × 1010 decays/s)
Answer #1

T1/2 =1.6*10^3 y

=1.6*10^3 y*(365.25 d/1y)*(24h/1d)*(60 min/1h)*(60 sec/1 min)

=5.049216*10^10 sec

decay constant (lamda) =(ln 2)/(T1/2) =1.37278*10^-11 /sec

decay rate =(2.6*10^13 atoms)*1.37278*10^-11 /sec

=356.9228 atoms/sec

decay rate =356.9228 atoms/sec*(1Ci /3.7*10^10 decays/sec)

=9.65*10^-9 Ci